OpenStudy (anonymous):

Solve: sin^(2) x - Sin x =0, for 0≤x≤3

OpenStudy (michele_laino):

here we have to change variable, namely: \[\Large \sin x = t\]

OpenStudy (michele_laino):

after that your equation can be rewritten as follows: \[\Large {t^2} - t = 0\]

OpenStudy (michele_laino):

please solve that equation for t

OpenStudy (anonymous):

So t=0?

OpenStudy (michele_laino):

hint, we can factor out t at left side, so we can write: \[\Large t\left( {t - 1} \right) = 0\]

OpenStudy (anonymous):

Oh, so t= 0 and 1?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now t=0 means: sin x=0 what is x?

OpenStudy (anonymous):

x= Radical 3 over 2?

OpenStudy (anonymous):

Oops!

OpenStudy (michele_laino):

no, the solution of that equation: sin x=0, are: \[x = 0,\pi ,2\pi ,3\pi ...\]

OpenStudy (michele_laino):

nevertheless only x=0, is acceptable, since pi >3

OpenStudy (anonymous):

Oh, i get it! Thank you!

OpenStudy (michele_laino):

please wait, we have to solve this equation: t=1 which means sin x =1

OpenStudy (michele_laino):

for which values x, we have: sin x= 1?

OpenStudy (michele_laino):

\[\sin x = 1 \Rightarrow x = ...?\]

OpenStudy (anonymous):

Pie over 2 and 3 pie over 2

OpenStudy (michele_laino):

no, the solutions are: \[x = \frac{\pi }{2},2\pi + \frac{\pi }{2},4\pi + \frac{\pi }{2}...\]

OpenStudy (anonymous):

Oh yea, my bad.

OpenStudy (michele_laino):

nevertheless only x= pi/2 is acceptable

OpenStudy (anonymous):

I get it. thanks :D

OpenStudy (michele_laino):

thanks! :)